# Nonlinear Model Order Reduction of Dynamical Systems in Process Engineering: Review and Comparison

### v1

@@ 3.2.  Non-intrusive general-purpose methods
Besides the classical intrusive reduction methods, recent advances in ML and system identification have opened new avenues toward data-driven non-intrusive model reduction [115, 116, 117]. Importantly, non-intrusive reduction does not operate on the FOM equations and is therefore “model-free”. In addition, many non-intrusive methods combine MOR and hyperreduction in a single step. However, notice that non-intrusive reduction may still exploit some structural knowledge, e.g., knowing that the FOM is a polynomial model [118]. Finally, the generic and homogeneous structure of non-intrusive ROMs can be exploited in optimization [119, 120]. As before, we disregard non-intrusive Krylov subspace methods [121, 122, 66], as these methods are not (yet) applicable to nonlinear multi-input systems.
@@ 3.2.1.  Dynamic mode decomposition
Among the conceptually most simple yet effective non-intrusive methods are dynamic mode decomposition (DMD) [19] and the eigensystem realization algorithm [123], both non-intrusive linear approaches for constructing linear ROMs, i.e., linear latent dynamics on a linear subspace. Both approaches employ SVD to identify and truncate the principal linear modes from trajectory snapshot data. An extension of DMD to systems with inputs or controls (DMDc) is presented in [124], providing a ROM:
@@
where $\mathbf{U}_{1}\in\mathbb{R}^{n_{x}\times n_{z}}$, $\mathbf{A}\in\mathbb{R}^{n_{z}\times n_{z}}$, and $\mathbf{B}\in\mathbb{R}^{n_{z}\times n_{u}}$. We refer to [124] for details on the computation of $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{U}_{1}$. For a stable linear FOM, [125] presents bounds on the prediction error, $\bm{e}(k)\triangleq\bm{x}_{\mathrm{r}}(k)-\bm{x}(k)$, for $k>N$. However, as discussed in [126], DMDc may struggle with significantly nonlinear systems.
@@ 3.2.2.  Applied Koopman theory
An extension of the DMD algorithm (EDMD) was proposed in [33]. Their approach still provides a linear ROM, but the mathematical procedure to construct the dynamics involves a prior nonlinear lifting step based on Koopman theory. Koopman theory [32, 127] postulates that a nonlinear dynamical system possesses a linear operator representation when lifted to a generally infinite-dimensional space of observables $g:\mathcal{X}\rightarrow\mathbb{C}$. Constructing a finite-dimensional or even low-order Koopman model by means of (E)DMD corresponds to a Galerkin projection of the Koopman operator onto a set of basis functions $g_{i}$, collected in $\bm{g}$. Hence, despite being a linear method, DMD is applicable to nonlinear systems, when substituting state snapshots $\bm{x}^{(k)}$ by a sufficiently rich set of embedded snapshots $\bm{g}(\bm{x}^{(k)})$. Common choices of real-valued observables include Hermite polynomials, radial basis functions, sparse regression, and ANNs [33, 128]. More recently, kernel methods have gained attention due to their strong approximation capabilities [129, 130, 131].
Notably, a Koopman model does not necessarily predict states but only deals with observables. However, we can adapt the Koopman framework to state prediction by demanding that $\bm{x}(t)$ can be inferred from $\bm{g}(\bm{x}(t))$. To this end, a common assumption is that $\bm{x}(t)$ can be linearly reconstructed from the nonlinear observables. For example, [132] and follow-ups [133, 134, 135] use EDMD with inputs or controls (EDMDc):
@@
In the past years, Koopman theory has become a popular approach for both model bilinearization [136, 137, 138] and non-intrusive MOR [139, 140]. For autonomous systems, [141] relaxes the assumption of a global linear projection and proposed to combine linear dynamics and nonlinear reconstruction by means of a decoder network $\bm{g}^{\dagger}$. In [51], we have shown that this strategy is also valid for input-affine systems. Under certain assumptions on the observables $\bm{g}$ (see [51]), the respective ROM takes the form:
@@
which is visualized by Fig. 3. Notice that the linear dynamics in Eq. 26 is a special case of the bilinear dynamics derived in [51]. For non-oscillating systems, we may choose a diagonal $\mathbf{A}$ and thereby reduce the trainable parameters. The combination of linear dynamics and nonlinear static map has “Wiener structure” and we term Eq. 26 a “Koopman-Wiener” (KW) model. In contrast to Eq. 19, we do not only project the nonlinear dynamics onto some subspace of $\mathcal{X}$, but we additionally demand the projection $\bm{g}$ to be linearizing. Consequently, the learning problem is more complex than standard MFL. Realizing the KW framework is most straightforward in a non-intrusive fashion, where an ML training combines reconstruction, prediction, and regularization loss terms [51].
@@
Finally, the flexibility of ML enables modifications of the KW structure. For example, the encoder may serve as a state estimator [142].
@@ 3.2.3.  Lift-and-project methods
Lift-and-project methods expand the FOM into an even higher-dimensional canonical form before projecting this representation to a low-dimensional subspace. The lifting step usually targets a linear, bilinear, or quadratic representation of the FOM. This extra step enables (i) the use of linear or bilinear MOR methods, and (ii) circumvents hyperreduction due to mathematically simpler ROMs. Methods such as Koopman lifting [143], Carleman bilinearization [144], piecewise linearization [145] or Taylor expansion [29, 30, 31], lay the theoretical foundation of lift-and-project methods.
Because intrusive variants of lift-and-project methods are mathematically involved and computationally expensive, non-intrusive implementations are convenient to circumvent lifting the FOM. For example, the studies [118, 146] presented a non-intrusive lifting method for polynomial FOMs, where snapshot data is lifted and POD-projected to learn a quadratic ROM. Therein, the model regression of the quadratic ROM is called “operator inference” (OPI) [147, 148]. However, this lift-and-learn method [118, 146] currently requires the manual (pen-and-paper) derivation of the candidate lifting maps.
@@ 3.2.4.  Reduced basis and manifold learning methods
An alternative non-intrusive reduction approach is to combine reduced basis methods or manifold learning with system identification methods. For example, given POD-reduced snapshots, the latent ROM dynamics may be learned using recurrent neural networks (RNNs) [149], temporal convolutional ANNs [150], or feedforward ANN predictors [151]. Analogously, given AEN-compressed snapshots, an ROM can be obtained, e.g., using RNNs [152] or Hammerstein-Wiener models [153]. These approaches are non-intrusive counterparts to Eq. 19.
The above methods are two-stage procedures because the slow manifold is constructed independently of the latent dynamics. However, this separation may introduce a bottleneck, i.e., the latent variables may not be suitable latent states, depending on the type of predictor. Interestingly, a similar bottleneck is also present in intrusive model reduction, as highlighted in [154, 104]. Non-intrusive single-stage approaches address this issue by learning slow manifold and latent dynamics simultaneously. Conveniently, the KW method reviewed above inherently follows this strategy [51]. In a similar but empirical setup, the works [155, 156] use AENs and linear time-invariant dynamics to learn stochastic ROMs. Alternative approaches combine MFL and RNNs [157], MFL and feedforward ANNs (termed MFL-ANN below) [158, 159, 160], or MFL and sparse identification of latent dynamics [161].
@@ 3.3.  Discussion and comparison
We present a theoretical comparison of the reviewed methods with respect to their nonlinear reduction properties in Table 1. First, we distinguish methods by their intrusive character (Intr). Intrusive MOR methods do not inherently perform hyperreduction (Hr) and thus require additional model simplification efforts for real-time applications. In contrast, non-intrusive methods use ROM structures that are typically (by design) much cheaper to evaluate so that a separate hyperreduction step is not needed.
Next, we distinguish whether the methods are snapshot-based (Snap). This categorization reveals another practical limitation (NU). Snapshot-based methods usually face the curse of dimensionality in $n_{u}$, since an independent excitation of inputs is commonly used to adequately cover the relevant state-space regions. As a consequence, snapshot-based methods are less straightforward to apply to systems with many inputs, i.e., $n_{u}\gg 1$. Hence, for systems with many independent inputs $\bm{u}(t)$, snapshot-free intrusive methods are the most suitable.
Despite the curse of dimensionality, input-response snapshots contain valuable information about controllability. As discussed in [162], when the snapshots are created via broad excitation of the system through the inputs, these data sample the controllable subspace. As a consequence, building a ROM based on input-response data inherently approximates the controllable subspace and eliminates the uncontrollable states (Ctr).
Table 1: Comparison of nonlinear MOR methods that preserve state information. Intr: intrusive method, Snap: snapshot-based method, Manif: type of submanifold, Lat: type of latent dynamics, Hr: hyperreduction features, Nu: suitable for many inputs ($n_{u}\gg 1$), Ctr: approximates controllable subspace, Stab: preserves stability. (✓) practical limitations.
@@
f) DMDc/EDMDc 3.2.1/2 Lin. Lin./Bilin. (✓)
@@
h) OPI 3.2.3 Lin. Quadr. (✓)
i) POD-ANN/RNN 3.2.4 Lin. Nonlin. (✓)
j) MFL-ANN/RNN 3.2.4 Nonlin. Nonlin. (✓)
@@
Lastly, we distinguish the reduction methods by the type of latent dynamics (Lat), see Eq. 4a. In general, the ROM of a nonlinear FOM is also nonlinear. However, DMDc and KW reduction methods construct linear latent dynamics, whereby the nonlinear ROM is simplified and “structurally regularized”. In other words, a ROM with linear latent dynamics cannot falsely exhibit strong nonlinear phenomena such as chaos or state multiplicity [163]. At the same time, these ROMs are limited to mildly nonlinear systems or short prediction horizons. However, a ROM with linear or bilinear latent dynamics is oftentimes sufficient, where more complex structures, e.g., MFL-ANN, bear a higher risk of overfitting and undesired dynamical behavior.
Relatively few nonlinear model reduction methods provide a-priori guarantees on stability preservation (Stab). These guarantees are established by system-theoretic arguments (SPT, AGG, POD-Res.) or through constrained learning problems (KW, OPI). Other methods require a careful a-posteriori investigation of stability properties of the ROM [77]. Regarding the ROM error, $\bm{e}(t)\triangleq\bm{x}_{\mathrm{r}}(t)-\bm{x}(t)$, available nonlinear MOR techniques usually do not provide universal error bounds but require a-posteriori ROM validation. Special MOR methods that provide error bounds for special types of nonlinear systems have been presented, e.g, in [164, 165].
@@ 4.  Reduced modeling of chemical processes
We now review specialized MOR methods developed in the PSE field as well as applications of general-purpose methods to PSE problems. In Section 4.1, we review three specializations of MOR methods from the previous section, namely (i) compartment modeling, (ii) model aggregation, and (iii) collocation method. Afterwards, in Section 4.2, we discuss the application of MOR to three very common unit operations: (i) distillation columns, (ii) heat exchangers (HX), and (iii) reactors. Again, we omit equations as far as possible and rather focus on an abstract discussion. Besides deterministic reduction, reduced modeling is often achieved based on knowledge or intuition through simplifying assumptions in the modeling process [166]. We also comment on these aspects.
@@ 4.1.1.  Compartment modeling
The key concept in compartment modeling (COMP) is the definition of interconnected functional subsystems, termed “compartments”, between the local microscale and the global system scale [167]. When applied for MOR, the introduction of compartments is often motivated by a time scale assumption, where local phenomena inside a compartment are assumed to be much faster than the overall response of the compartment [85]. Given a discretized model of a conservative distributed system, a compartment is a cluster of neighboring cells, and the compartment equations are obtained by superposition of the extensive balance equations, e.g., absolute enthalpy or total molar holdup, of these cells. In terms of intensive quantities, e.g., molar enthalpy and molar fraction, this superposition corresponds to a mass-weighted averaging.
By applying the above time scale assumption, the local dynamics of cells within a compartment are degenerated to quasi-stationary equations, and we obtain a ROM. In addition, replacing quasi-stationary equations by shortcut methods, averaging rules, or regression models realizes a hyperreduction. COMP is a heuristic approach to transform an FOM into Eqs. 12 and 13. Hence, we regard COMP as a heuristic SPT method, inheriting properties such as the potential appearance of a non-physical IR (see discussion below).

### v2

@@ 3.2. Non-intrusive general-purpose methods
Besides the classical intrusive reduction methods, recent advances in machine learning and system identification have opened new avenues toward data-driven non-intrusive model reduction [142, 34, 81]. Importantly, non-intrusive reduction does not operate on the FOM equations and is therefore model-free. In addition, many non-intrusive methods combine MOR and hyperreduction in a single step. However, notice that non-intrusive reduction may still exploit some structural knowledge, e.g., knowing that the FOM is a polynomial model [149]. Finally, the generic and homogeneous structure of non-intrusive ROMs can be exploited in optimization [263, 260]. Like Krylov subspace methods, we disregard non-intrusive rational approximation [189, 218, 253, 198], as these methods are not (yet) applicable to nonlinear multi-input systems.
@@ 3.2.1. Dynamic mode decomposition
Among the conceptually most simple yet effective non-intrusive methods is dynamic mode decomposition (DMD) [257, 103], a non-intrusive linear approach to construct linear ROMs, i.e., linear latent dynamics on a linear subspace. DMD employs SVD to identify and truncate the principal linear modes from trajectory snapshot data. An extension of DMD to systems with inputs or controls (DMDc) is presented in [229], providing a ROM:
@@
where $\mathbf{U}_{1}\in\mathbb{R}^{n_{x}\times n_{z}}$, $\mathbf{A}\in\mathbb{R}^{n_{z}\times n_{z}}$, and $\mathbf{B}\in\mathbb{R}^{n_{z}\times n_{u}}$. We refer to [229] for details on the computation of $\mathbf{A}$, $\mathbf{B}$, and $\mathbf{U}_{1}$. For a stable linear FOM, [171] presents bounds on the prediction error, $\bm{e}(k)\triangleq\bm{x}_{\mathrm{r}}(k)-\bm{x}(k)$, for $k>N$. However, as discussed in [129], DMDc may encounter inaccuracies when applied to significantly nonlinear systems.
@@ 3.2.2. Applied Koopman theory
An extension of the DMD algorithm was proposed in [298]. Their approach still provides a linear ROM, but the mathematical procedure to construct the dynamics involves a prior nonlinear lifting step based on Koopman theory. Koopman theory [191, 212] postulates that a nonlinear dynamical system possesses a linear operator representation when lifted to a generally infinite-dimensional space of observables $g:\mathcal{X}\rightarrow\mathbb{C}$. Constructing a finite-dimensional or even low-order Koopman model by means of (E)DMD corresponds to a Galerkin projection of the Koopman operator onto a set of basis functions $g_{i}$, collected in a vector $\bm{g}$. Hence, despite being a linear method, DMD is applicable to nonlinear systems, when substituting state snapshots $\bm{x}^{(k)}$ by a sufficiently rich set of embedded snapshots $\bm{g}(\bm{x}^{(k)})$. Common choices of real-valued observables include Hermite polynomials, radial basis functions, sparse regression, and ANNs [298, 165]. More recently, kernel methods have gained attention due to their strong approximation capabilities [56, 112, 143].
Notably, a Koopman model does not necessarily predict states but only deals with observables. However, we can adapt the Koopman framework to state prediction by demanding that $\bm{x}(t)$ can be inferred from $\bm{g}(\bm{x}(t))$. To this end, a common assumption is that $\bm{x}(t)$ can be linearly reconstructed from the nonlinear observables. For example, [146] and follow-ups [199, 108, 107] use extended DMDc:
@@
In the past years, Koopman theory has become a popular approach for both model bilinearization [297, 276, 221] and non-intrusive MOR [211, 220]. For autonomous systems, [174] relaxes the assumption of a global linear projection and proposed to combine linear dynamics and nonlinear reconstruction by means of a decoder network $\bm{g}^{\dagger}$. In [261], we have shown that this strategy is also valid for input-affine systems. Under certain assumptions on the observables $\bm{g}$, see [261], the respective ROM takes the form:
@@
which is visualized by Fig. 2. Notice that the linear dynamics in (27) is a special case of the bilinear dynamics derived in [261]. For non-oscillating systems, we may choose a diagonal $\mathbf{A}$ and thereby reduce the trainable parameters. The combination of linear dynamics and nonlinear static map has Wiener structure and we term (27) a Koopman-Wiener (KW) model. In contrast to (20), we do not only project the nonlinear dynamics onto some subspace of $\mathcal{X}$, but we additionally demand the projection $\bm{g}$ to be linearizing. Consequently, the learning problem is more complex than standard MFL. Realizing the KW framework is most straightforward in a non-intrusive fashion, where a training problem combines reconstruction, prediction, and regularization loss terms [261].
@@
Finally, the flexibility of machine learning enables modifications of the KW structure. For example, the encoder may serve as a state estimator [262].
@@ 3.2.3. Lift-and-project methods
Lift-and-project methods expand the FOM into an even higher-dimensional canonical form before projecting this representation to a low-dimensional subspace. The lifting step usually targets a linear, bilinear, or quadratic representation of the FOM. Notably, this extra step enables (i) the use of linear or quadratic/polynomial MOR methods, and (ii) circumvents hyperreduction due to mathematically simpler ROMs. Methods such as Koopman lifting [147], Carleman bilinearization [273], piecewise linearization [237], variational analysis via Taylor expansion [226, 66], and quadratization [95, 115, 39] lay the theoretical foundation of lift-and-project methods. A detailed overview is given in [18] and software to automate the lifting was recently presented in [39, 116, 65].
In cases where intrusive variants of lift-and-project methods are mathematically involved or too computationally expensive, non-intrusive implementations are convenient to circumvent lifting the FOM. For example, the studies [149, 230] presented a non-intrusive lifting method for polynomial FOMs, where snapshot data is lifted and POD-projected to learn a quadratic ROM. Therein, the model regression of the quadratic ROM is called operator inference [219, 21]. Extensions of operator inference to polynomial manifolds and ROMs are found, e.g., in [78, 77]. Also, the non-intrusive (extended) DMDc and KW methods can be regarded as a variant of operator inference, where the latent dynamics are restricted to a linear or bilinear form.
@@ 3.2.4. Reduced basis and manifold learning methods
An alternative non-intrusive reduction approach is to combine reduced basis methods or manifold learning with system identification methods. For example, given POD-reduced snapshots, the latent ROM dynamics may be learned using recurrent neural networks [291], temporal convolutional ANNs [300], or feedforward ANN predictors [76]. Analogously, given AEN-compressed snapshots, an ROM can be obtained, e.g., using ANNs [295] or Hammerstein-Wiener models [283]. These approaches are non-intrusive counterparts to (20). Non-intrusive operator inference for quadratic manifolds was presented in [78], where the authors infer a quadratic ROM from POD-projected snapshot data. This approach was extended to polynomial manifolds and ROMs in [77].
The above methods are two-stage procedures because the slow manifold is constructed independently of the latent dynamics. However, this separation may introduce a bottleneck, i.e., the latent variables may not be suitable latent states, depending on the type of predictor. Interestingly, a similar bottleneck is also present in intrusive model reduction, as highlighted in [210, 209]. Non-intrusive single-stage approaches address this issue by learning slow manifold and latent dynamics simultaneously. Conveniently, the KW method reviewed above inherently follows this strategy [261]. In a similar but empirical setup, the works [292, 89] use AENs and linear time-invariant dynamics to learn stochastic ROMs. A combination of AEN and quadratic dynamics is used in [91]. Alternative approaches combine MFL and recurrent ANNs [86] or feedforward ANNs [8, 187, 73, 85] (termed MFL-ANN below), as well as MFL and sparse identification of latent dynamics [53].
@@ 3.3. Discussion and comparison
We present a comparison of the reviewed methods with respect to their nonlinear reduction properties in Table 2. First, we distinguish methods by their intrusive character (Intr). Intrusive MOR methods do not inherently perform hyperreduction (Hr) and thus require additional model simplification efforts for real-time applications. In contrast, non-intrusive methods use ROM structures that are typically (by design) much cheaper to evaluate so that a separate hyperreduction step is not needed.
Next, we distinguish whether the methods are snapshot-based (Snap). This categorization reveals another practical limitation, namely the ability to deal with a large number of inputs (Nu). Snapshot-based methods usually face the curse of dimensionality in $n_{u}$, since an independent excitation of inputs is commonly used to adequately cover the relevant state-space regions. As a consequence, snapshot-based methods are less straightforward to apply to systems with many inputs, i.e., $n_{u}\gg 1$. Hence, for systems with many independent inputs $\bm{u}(t)$, snapshot-free intrusive methods are the most suitable.
Despite the curse of dimensionality in data sampling, input-response snapshots contain valuable information about controllability. As discussed in [296], when the snapshots are created via broad excitation of the system through the inputs, these data sample the controllable subspace. As a consequence, building a ROM based on input-response data inherently approximates the controllable subspace and eliminates the uncontrollable states (Ctr).
Table 2: Comparison of nonlinear MOR methods that preserve state information. Intr: intrusive method, Snap: snapshot-based method, Manif: type of submanifold, Lat: type of latent dynamics, Hr: hyperreduction features, Nu: suitable for many inputs ($n_{u}\gg 1$), Ctr: approximates controllable subspace, Stab: preserves stability. ✓: yes, (✓): practical limitations, ✗: no.
@@
f) (Extended) DMDc 3.2.1/2 Lin. Lin./Bilin. (✓)
@@
h) Operator inference 3.2.3 Polyn. Polyn. (✓)
i) POD-ANN 3.2.4 Lin. Nonlin. (✓)
j) MFL-ANN 3.2.4 Nonlin. Nonlin. (✓)
@@
Lastly, we distinguish the reduction methods by the type of latent dynamics (Lat), see (4a). In general, the ROM of a nonlinear FOM is also nonlinear. However, DMDc and KW reduction methods construct linear latent dynamics, whereby the nonlinear ROM is simplified and structurally regularized. In other words, a ROM with linear latent dynamics cannot falsely exhibit strong nonlinear phenomena such as chaos or state multiplicity [217]. At the same time, these ROMs are limited to mildly nonlinear systems or short prediction horizons. However, a ROM with linear or bilinear latent dynamics is oftentimes sufficient, where more complex structures, e.g., MFL-ANN, bear a higher risk of overfitting and undesired dynamical behavior.
Only a subgroup of nonlinear model reduction methods provide a-priori guarantees on stability preservation (Stab). These guarantees are established by system-theoretic arguments (SPT, AGG, POD-Res.) or through constrained learning problems (DMD, KW, and operator inference). In reduced basis methods, stability preservation may be enabled through variable transformation or modification of the inner product [248, 16]. Other methods require a careful a-posteriori investigation of ROM stability [22, 117]. We also mention methods developed to incorporate physics priors, particularly for Hamiltonian [2, 266, 94, 36], port-Hamiltonian [96, 190, 79], and Lagrangian systems [42, 75]. These formalisms support the construction of ROMs that respect energy and mass conservation and preserve stability [36, 162]. Here, the port-Hamiltonian framework is particularly suitable for dynamical systems with inputs or interconnected process subsystems.
Regarding the ROM error, $\bm{e}(t)\triangleq\bm{x}_{\mathrm{r}}(t)-\bm{x}(t)$, many publications on MOR techniques do not derive universal error bounds but perform a-posteriori ROM validation. Rigorous error bounds are presented, e.g., in [10, 140, 26, 265, 138]. The accuracy of a ROM is directly coupled to modal closure and memory effects [3]. Hence, methods for ROM closure modeling, e.g., the Mori-Zwanzig formalism [90], support tighter error bounds.
@@ 4. Reduced modeling of chemical processes
We now review specialized MOR methods developed in the process systems engineering field as well as applications of general-purpose methods to such problems. In Section 4.1, we review three specializations of MOR methods from the previous section, namely (i) compartment modeling, (ii) model aggregation, and (iii) collocation method. Afterwards, in Section 4.2, we discuss the application of MOR to three very common unit operations: (i) distillation columns, (ii) heat exchangers (HX), and (iii) reactors. Again, we omit equations as far as possible and rather focus on an informal discussion. Besides deterministic reduction, reduced modeling is often achieved based on knowledge or intuition through simplifying assumptions in the modeling process [64]. We also comment on these aspects.
@@ 4.1.1. Compartment modeling
The key concept in compartment modeling (COMP) is the definition of interconnected functional subsystems, termed compartments, between the local microscale and the global system scale [134]. When applied for MOR, the introduction of compartments is often motivated by a time scale assumption, where local phenomena inside a compartment are assumed to be much faster than the overall response of the compartment [20]. Given a discretized model of a conservative distributed system, a compartment is a cluster of neighboring cells, and the compartment equations are obtained by superposition of the extensive balance equations, e.g., absolute enthalpy or total molar holdup, of these cells. In terms of intensive quantities, e.g., molar enthalpy and molar fraction, this superposition corresponds to a mass-weighted averaging.
By applying the above time scale assumption, the local dynamics of cells within a compartment are degenerated to quasi-stationary equations, and we obtain a ROM. In addition, replacing quasi-stationary equations by shortcut methods, averaging rules, or regression models realizes a hyperreduction. COMP is a heuristic approach to transform an FOM into Eqs. (13) and (14). Hence, we regard COMP as a heuristic SPT method, inheriting properties such as the potential appearance of a non-physical inverse response.